# RemixFusion: Residual-based Mixed Representation for Large-scale Online RGB-D Reconstruction

### v1

@@ Abstract.
The introduction of the neural implicit representation has notably propelled the advancement of online dense reconstruction techniques. Compared to traditional explicit representations, such as TSDF, it substantially improves the mapping completeness and memory efficiency. However, the lack of reconstruction details and the time-consuming learning of neural representations hinder the widespread application of neural-based methods to large-scale online reconstruction. We introduce RemixFusion, a novel residual-based mixed representation for scene reconstruction and camera pose estimation dedicated to high-quality and large-scale online RGB-D reconstruction. In particular, we propose a residual-based map representation comprised of an explicit coarse TSDF grid and an implicit neural module that produces residuals representing fine-grained details to be added to the coarse grid. Such mixed representation allows for detail-rich reconstruction with bounded time and memory budget, contrasting with the overly-smoothed results by the purely implicit representations, thus paving the way for high-quality camera tracking. Furthermore, we extend the residual-based representation to handle multi-frame joint pose optimization via bundle adjustment (BA). In contrast to the existing methods, which optimize poses directly, we opt to optimize pose changes. Combined with a novel technique for adaptive gradient amplification, our method attains better optimization convergence and global optimality. Furthermore, we adopt a local moving volume to factorize the whole mixed scene representation with a divide-and-conquer design to facilitate efficient online learning in our residual-based framework. Extensive experiments demonstrate that our method surpasses all state-of-the-art ones, including those based either on explicit or implicit representations, in terms of the accuracy of both mapping and tracking on large-scale scenes. The code will be released at project page.
@@ 1. INTRODUCTION
Online dense reconstruction based on RGB-D cameras has made significant advances in recent years and lately, the progress has been propelled by the fast development of robust tracking based on randomized optimization (Zhang et al., 2021; Tang et al., 2023; Zhang et al., 2022) and scalable mapping based on neural map representation (Wang et al., 2021; Chen et al., 2022; Li et al., 2023; Wang et al., 2022; Azinović et al., 2022). The latter has been incurring an in-depth revolution in scene representation methods, from explicit volumetric fields or point clouds to implicit neural fields of occupancy or radiance.
Implicit scene representation leads to improved completeness of dense reconstructions, yet it also introduces new challenges in scene mapping and pose tracking. These challenges are particularly evident in large-scale scene reconstruction. As the size of the scenes scales up, it becomes increasingly difficult to balance the high-precision reconstruction of details with the significant computation and memory overheads required by online systems. In the traditional volumetric fusion approach (Izadi et al., 2011), depth maps are directly fused into a TSDF field. The reconstructed details can be aligned with the depth maps as long as sufficient TSDF resolution is adopted albeit at the expense of higher storage cost. When working with implicit scene representations (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023; Johari et al., 2023; Zhang et al., 2023), however, this issue is not straightforward to address as reconstruction details are encoded with neural networks. Even with larger networks and extended training time, encoding high-frequency geometric information remains difficult, and the problem is exacerbated as the scene scales up. Consequently, to make an online dense reconstruction system real-time capable with a limited memory footprint, neither explicit nor implicit scene representation can balance between effective modeling of fine-grained details and efficient reconstruction of scenes at scale.
@@
Residual learning has been proven in various works (He et al., 2016; Xiangli et al., 2022; Kim et al., 2016) to improve the convergence of neural network training, enhance generality, and enable the network to encode rich details. These advantages are especially useful when learning implicit representations for large-scale scene reconstruction. Motivated by that, we propose RemixFusion, adopting the concept of residuals with a mixed representation of implicit and explicit, for both detail-rich mapping and accurate pose tracking in online dense reconstruction. To address the inefficiency of implicit representations in capturing high-frequency details in large scenes, we first combine explicit and implicit scene representations in a residual form. The base of our mixed representation is a coarse-grained explicit 3D grid volume that stores low-frequency scene structure. High-frequency geometry details are captured in the parameters of the neural module, coupling with the coarse-grained base as a residual. This approach reduces memory overhead by lowering the resolution of the explicit representation while preserving scene details in a memory-efficient form through neural representation. Furthermore, the training complexity of the residual neural module is significantly reduced since it focuses on encoding only high-frequency features. As a result, this mixed representation not only preserves fine-grained details in a memory-efficient way but also greatly improves online learning efficiency.
@@
For pose estimation, based on the mixed representation, we propose a residual-based multi-frame pose optimization method. Specifically, we estimate residual pose corrections to refine the initial poses from front-end tracking, thereby enhancing geometric consistency during bundle adjustment (BA). By encoding only residual poses in the network, our approach allows the BA based on implicit representations to focus on optimizing pose changes rather than absolute poses, thus improving the learning efficiency and multi-view consistency. In practice, the real-time constraints make it impractical to perform bundle adjustment on the whole reconstruction. The sparse sampling of the scene results in discontinuous perception of surface details during joint optimization, which in turn reduces the optimization’s ability to escape local minima. To address this challenge, we introduce an adaptive gradient amplification based on the reconstructed surface, allowing the BA to obtain better optimization gradients even when the detail perception is discontinuous in real-world large-scale scenarios.
@@
As shown in Figure 1, our online dense reconstruction method, employing an effective fusion of explicit and implicit representations, achieves fine-grained reconstruction of large-scale scenes with a relatively low GPU memory cost. Our method exhibits significantly improved online learning efficiency thanks to the explicit map, resulting in a frame rate that is 2.5 times higher than that of other methods based solely on implicit representation. Furthermore, the residual-based camera pose bundle adjustment benefits pose tracking, which improves tracking accuracy by over the state-of-the-art dense SLAM (Johari et al., 2023) on BS3D (Mustaniemi et al., 2023). In summary, our contributions include:
@@
We propose a residual-based bundle adjustment technique that employs a tiny MLP for residual-based pose refinement. Compared to traditional BAs, our method improves pose estimation in terms of both efficiency and robustness.
We have implemented an efficient system of online RGB-D dense reconstruction which realizes robust and fine-grained real-time reconstruction for large scenes over 1000 with an affordable GPU memory footprint.
@@ 2. RELATED WORK
The explicit scene representation has been widely studied in the last decades, most including voxel-based and point-based methods. In the context of voxel-based methods, KinectFusion (Izadi et al., 2011) proposes the Truncated Signed Distance Function (TSDF) for encoding the scenes with a consumer depth sensor. Considering the waste of encoding the empty space of the scene, follow-up works (Whelan et al., 2012; Roth and Vona, 2012) extend the KinectFusion by leveraging a moving TSDF volume to encode larger scenes. (Nießner et al., 2013a; Dai et al., 2017b) encode the scenes with hash blocks to further improve the scalability of scene reconstruction. Global bundle adjustment is utilized to reduce the drift in pose estimation and enhance model consistency in terms of large-scale indoor environments (Dai et al., 2017c). For points-based methods, (Keller et al., 2013; Whelan et al., 2015, 2016) leverages points and surfels to achieve scalability and flexibility for encoding scenes. With explicit scene representation, these methods leverage classic second-order optimization methods, such as Gaussian-Newton or Levenberg-Marquardt (LM). Recently, (Zhang et al., 2021, 2022) propose random optimization to improve the robustness of camera tracking and demonstrate good performance on large-scale fast-moving motions. However, explicit methods suffer from memory consumption, especially in large-scale scenes.
@@
Reconstruction of large-scale scenes requires efficient representations. As for explicit voxel-based methods, hashing schemes (Nießner et al., 2013b; Dai et al., 2017c) and octrees (Liu et al., 2020; Mao et al., 2023) are adopted to reduce the memory footprint. However, these methods do not reduce the number of parameters in modeling and lack the ability of completion. In terms of implicit methods, learnable hash grids (Wang et al., 2023; Zhang et al., 2023), tri-planes (Johari et al., 2023) and neural points (Sandström et al., 2023; Hu et al., 2023) are utilized to achieve the trade-off of efficiency and reconstruction quality. However, these methods suffer from the high latency of the model update. Another line of work leverages submaps (Tancik et al., 2022; Tang et al., 2023; Mao et al., 2023) to dynamically allocate the implicit maps with the moving camera. The multi-submap techniques improve the scalability while costing much more time in submap management. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) shows promising results in novel view synthesis. A series of works (Keetha et al., 2024; Huang et al., 2024a; Matsuki et al., 2024; Ha et al., 2024) explore applying 3DGS to monocular or RGB-D SLAM. Submaps and loop closure techniques (Zhu et al., 2025) are incorporated to improve the tracking accuracy. Furthermore, a more compact representation (Peng et al., 2024) is proposed to enhance the efficiency of 3DGS. These methods demonstrate high-fidelity rendering and generalization capabilities in different scenarios. However, these 3DGS-based methods are still less memory-friendly than implicit methods since they need numerous 3DGS for the explicit model, especially in large-scale scenes. We propose to take advantage of both explicit and implicit representations to remain both memory-efficient and time-efficient via residual mixing.
@@ 3.1. Overview
In this section, we first introduce our residual-based framework, including the explicit-implicit mixed scene representation and residual-based bundle adjustment for camera poses. Our insight behind these designs is that the residual can be adopted to force networks to focus on the high-frequency details upon existing coarse geometry, relieving the burden of neural learning and enhancing efficiency. In this way, the proposed online dense reconstruction framework is both storage-efficient and time-efficient, which is crucial for application in large-scale scenes with real-time requirements.
More specifically, as shown in Eq. 1, we decouple the complete representation into a coarse representation and a residual one , which are optimized during reconstruction and pose estimation. In general, is expected to be fast and robust, such as TSDF, providing the coarse geometry seamlessly with the input data. is supposed to be expressive and capture the fine-grained details based on . The coarse estimation is usually constructed based on explicit representations, and the neural networks are adopted for residual refinement . The coarse estimation and residual refinement are performed alternately for both reconstruction and pose estimation.
@@
The overview of our method is demonstrated in Figure 2. Given the successive RGB-D inputs, we first leverage the scalable randomized optimization with limited and fixed footprints for the coarse pose estimation (Figure 2(a)). Keyframes are selected from these input frames. The corresponding initial camera poses are fed into a tiny MLP for the joint residual pose refinement (Figure 2(b)).
Note that the proposed residual-based BA of the camera poses is optimized by the gradients given by the multi-view constraints based on the reconstructed scene. To support pose estimation better, an explicit-implicit scene representation mixture is introduced simultaneously. The online learning (Figure 2(c)) of this representation depends on the optimized camera poses. Specifically, we perform ray casting and sample many points along the ray directions accordingly. Embeddings of these points, including the hash and explicit embedding together with positional embedding, are sent to two MLPs for the TSDF and RGB residual prediction. The residuals indicating the high-frequency details are added to the explicit coarse bases, outputting the final reconstruction. The residual-based BA and residual-based mapping are running alternately with continuously updated global reconstruction and optimized keyframe poses.
@@ 3.2. Residual-based Mapping
In large-scale scenes, various layouts and complex structures make it challenging for the scene representations to remain both expressive and efficient. Shown in Figure 2(c), we employ a mixed representation to address this challenge. This representation consists of a explicit coarse TSDF grid and an implicit neural map , which correspond to and in Eq. 1 respectively. The key insight of the residual-based representations is to obtain the coarse reconstruction in real time and provide the basis surface for the implicit module to efficiently deform and refine. Consequently, given an arbitrary 3D position , the attributes which are the geometry and appearance of can be formed in a residual-based aggregation as:
@@
Given the posed RGB-D frames, the explicit component of the proposed scene representation is constructed via TSDF Fusion (Curless and Levoy, 1996) with a relatively low resolution to store the coarse TSDF and RGB attributes, serving as explicit bases of our reconstruction. Similar to some previous implicit-based methods, the implicit component includes two kinds of embedding functions, which are hash embedding (Müller et al., 2022) and positional embedding (Müller et al., 2019), to encode the reconstruction residuals for an arbitrary position upon the . Specifically, adopting this joint embedding in the representation can balance the training efficiency and reconstruction quality(Wang et al., 2023), which is critical for online reconstruction. Furthermore, the residual learning can not be independent without the explicit bases , the final residual-based aggregation to calculate the attributes for a position can be formed based on Eq. 2 as:
@@
Different from the can be reconstructed directly based on the fusion of the posed RGB-D frames, and in our implicit neural map need to be carefully optimized during the reconstruction ( is a constant projection). Therefore, proper supervision for learning and formulating a good loss function is critical here.
Following (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023), we adopt the input RGB-D frame sequence with the estimated poses as the approximation of ground-truth modeling to provide the supervision in a projection-and-measure fashion. Volume rendering is leveraged to output the re-projected RGB-D images based on the weights of points on the casting ray from each image pixel. Similar to (Azinović et al., 2022), the weights are obtained by the Sigmoid functions and as shown in Eq. 5.
Given an image pixel and the corresponding sampled points on its casting rays, the rendered RGB values and depth values can be obtained as Eq. 6 and Eq. 7.
where is the 3D position of under the camera coordinates with the corresponding pose and denotes truncation threshold adopted in the TSDF construction.
One more thing, due to the resolution of being low, the truncation threshold for it needs to be larger than the one adopted in the residual component . To align these two modules, the explicit bases adopted in TSDF component of Eq. 3 needs to be specified as:
The is the clamp function which ensures to be in the range of with threshold . The aligned explicit bases is only adopted for the TSDF calculation, while with the direct trilinear interpolation is proper for the RGB prediction.
@@
For the running efficiency, we sample pixels in each iteration for neural learning. More details about sampling strategies can be found in Section 3.4. We use the observed RGB and depth images for supervision. The photometric and geometric loss are as follows:
@@
where is the random sampled vertices on the hash grids and means the difference of hash features between the adjacent vertices along these three axes.

### v3

@@ Abstract.
The introduction of the neural implicit representation has notably propelled the advancement of online dense reconstruction techniques. Compared to traditional explicit representations, such as TSDF, it substantially improves the mapping completeness and memory efficiency. However, the lack of reconstruction details and the time-consuming learning of neural representations hinder the widespread application of neural-based methods to large-scale online reconstruction. We introduce RemixFusion, a novel residual-based mixed representation for scene reconstruction and camera pose estimation dedicated to high-quality and large-scale online RGB-D reconstruction. In particular, we propose a residual-based map representation comprised of an explicit coarse TSDF grid and an implicit neural module that produces residuals representing fine-grained details to be added to the coarse grid. Such mixed representation allows for detail-rich reconstruction with bounded time and memory budget, contrasting with the overly-smoothed results by the purely implicit representations, thus paving the way for high-quality camera tracking. Furthermore, we extend the residual-based representation to handle multi-frame joint pose optimization via bundle adjustment (BA). In contrast to the existing methods, which optimize poses directly, we opt to optimize pose changes. Combined with a novel technique for adaptive gradient amplification, our method attains better optimization convergence and global optimality. Furthermore, we adopt a local moving volume to factorize the whole mixed scene representation with a divide-and-conquer design to facilitate efficient online learning in our residual-based framework. Extensive experiments demonstrate that our method surpasses all state-of-the-art ones, including those based either on explicit or implicit representations, in terms of the accuracy of both mapping and tracking on large-scale scenes.
@@ 1. INTRODUCTION
Online dense reconstruction based on RGB-D cameras has made significant advances in recent years. Lately, the progress has been propelled by the fast development of robust tracking based on randomized optimization (Zhang et al., 2021, 2022) and scalable mapping based on neural map representation (Wang et al., 2021; Chen et al., 2022; Li et al., 2023). The latter has been incurring an in-depth revolution in scene representation methods, from explicit volumetric fields or point clouds to implicit neural fields of occupancy or radiance.
Implicit scene representation leads to improved completeness of dense reconstructions, yet it also introduces new challenges in scene mapping and pose tracking. These challenges are particularly evident in large-scale scene reconstruction. As the size of the scenes scales up, it becomes increasingly difficult to balance the high-precision reconstruction of details with the significant computation and memory overheads required by online systems. In the traditional volumetric fusion approach (Izadi et al., 2011), depth maps are directly fused into a TSDF field. The reconstructed details can be aligned with the depth maps as long as sufficient TSDF resolution is adopted, albeit at the expense of higher storage cost. When working with implicit scene representations (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023; Johari et al., 2023; Zhang et al., 2023), however, this issue is not straightforward to address as reconstruction details are encoded with neural networks. Even with larger networks and extended training time, encoding high-frequency geometric information remains difficult, and the problem is exacerbated as the scene scales up. Consequently, to make an online dense reconstruction system real-time capable with a limited memory footprint, neither explicit nor implicit scene representation can balance between effective modeling of fine-grained details and efficient reconstruction of scenes at scale.
@@
Residual learning has been proven in various works (He et al., 2016; Xiangli et al., 2022; Kim et al., 2016; Qin et al., 2022) to improve the convergence of neural network training, enhance generality, and enable the network to encode rich details. These advantages are especially useful when learning implicit representations for large-scale scene reconstruction. Motivated by that, we propose RemixFusion, adopting the concept of residuals with a mixed representation of implicit and explicit, for both detail-rich mapping and accurate pose tracking in online dense reconstruction. To address the inefficiency of implicit representations in capturing high-frequency details in large scenes, we first combine explicit and implicit scene representations in a residual form. The base of our mixed representation is a coarse-grained explicit 3D grid volume that stores low-frequency scene structure. High-frequency geometry details, which serve as residuals, are captured in the parameters of the neural module and coupled with the coarse-grained base. This approach reduces memory overhead by lowering the resolution of the explicit representation while preserving scene details in a memory-efficient form through neural representation. Furthermore, the training complexity of the residual neural module is significantly reduced since it focuses on encoding only high-frequency features. Thus, this mixed representation not only preserves fine-grained details in a memory-efficient way but also improves online learning efficiency.
@@
For pose estimation, based on the mixed representation, we propose a residual-based multi-frame pose optimization method. Specifically, we estimate residual pose corrections to refine the initial poses from front-end tracking, thereby enhancing geometric consistency during bundle adjustment (BA). By encoding only residual poses in the network, our approach allows the BA based on implicit representations to focus on optimizing pose changes rather than absolute poses, thus improving the learning efficiency and multi-view consistency. In practice, the real-time constraints make it impractical to perform bundle adjustment on the whole reconstruction. The sparse sampling of the scene results in discontinuous perception of surface details during joint optimization, which in turn reduces the optimization’s ability to escape local minima. To address this challenge, we introduce an adaptive gradient amplification based on the reconstructed surface, allowing the BA to obtain better optimization gradients even when the detailed perception is discontinuous in real-world large-scale scenarios.
@@
As shown in Figure 1, our online dense reconstruction method, employing an effective fusion of explicit and implicit representations, achieves fine-grained reconstruction of large-scale scenes with a relatively low GPU memory cost. Our method exhibits significantly improved online learning efficiency thanks to the explicit map, resulting in a frame rate that is 2.4 times higher than that of other methods based solely on implicit representation. Furthermore, the residual-based camera pose bundle adjustment benefits pose tracking, which improves tracking accuracy by over the state-of-the-art dense SLAM (Johari et al., 2023) on BS3D (Mustaniemi et al., 2023). In summary, our contributions include:
@@
We propose a residual-based bundle adjustment technique that employs a tiny MLP for residual-based pose refinement. Compared to traditional BA, our method improves pose estimation in terms of both efficiency and robustness.
We have implemented an efficient system of online RGB-D dense reconstruction, which realizes robust and fine-grained real-time reconstruction for large scenes over 1000 with an affordable GPU memory footprint.
@@ 2. RELATED WORK
The explicit representation has been widely studied in the last decades, including voxel-based and point-based methods. In the context of voxel-based methods, KinectFusion (Izadi et al., 2011) proposes the Truncated Signed Distance Function (TSDF) for encoding the scenes with a consumer depth sensor. Considering the inefficiency of encoding the empty space of the scene, follow-up works (Whelan et al., 2012; Roth and Vona, 2012) extend the KinectFusion by leveraging a moving TSDF volume to encode larger scenes. (Nießner et al., 2013a; Dai et al., 2017b) encode the scenes with hash blocks to further improve the scalability of scene reconstruction. Global bundle adjustment is utilized to reduce the drift in pose estimation and enhance model consistency in terms of large-scale indoor environments (Dai et al., 2017c). For points-based methods, (Keller et al., 2013; Whelan et al., 2015, 2016) leverages points and surfels to achieve scalability and flexibility for encoding scenes. With explicit scene representation, these methods leverage classic second-order optimization methods, such as Gaussian-Newton or Levenberg-Marquardt (LM). Recently, (Zhang et al., 2021, 2022) propose random optimization to improve the robustness of camera tracking and demonstrate good performance on large-scale fast-moving motions. However, explicit methods suffer from memory consumption, especially in large-scale scenes.
@@
Reconstruction of large-scale scenes requires efficient representations. As for explicit voxel-based methods, hashing schemes (Nießner et al., 2013b; Dai et al., 2017c) and octrees (Liu et al., 2020; Mao et al., 2023) are adopted to reduce the memory footprint. However, these methods only reduce the storage required for empty space and lack the ability of completion. In terms of implicit methods, learnable hash grids (Wang et al., 2023; Zhang et al., 2023), tri-planes (Johari et al., 2023), and neural points (Sandström et al., 2023; Hu et al., 2023) are utilized to achieve the trade-off of efficiency and reconstruction quality. However, these methods suffer from the high latency of the model update. Another line of work leverages submaps (Tancik et al., 2022; Tang et al., 2023; Mao et al., 2023) to dynamically allocate the implicit maps with the moving camera. The multi-submap techniques improve the scalability while costing much more time in submap management. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) shows promising results in novel view synthesis. A series of works (Keetha et al., 2024; Huang et al., 2024a; Matsuki et al., 2024; Ha et al., 2024) explore applying 3DGS to monocular or RGB-D SLAM. Submaps and loop closure techniques (Zhu et al., 2025) are incorporated to improve the tracking accuracy. Furthermore, a more compact representation (Peng et al., 2024) is proposed to enhance the efficiency of 3DGS. These methods demonstrate high-fidelity rendering and generalization capabilities in different scenarios. However, these 3DGS-based methods are still less memory-friendly than implicit methods since they need numerous 3DGS for the explicit model, especially in large-scale scenes. We propose to take advantage of both explicit and implicit representations to remain both memory-efficient and time-efficient via residual mixing.
@@ 3.1. Overview
In this section, we first introduce our residual-based framework, including the explicit-implicit mixed scene representation and residual-based bundle adjustment for camera poses. Our insight behind these designs is that the residual can be adopted to force networks to focus on the high-frequency details upon the existing coarse geometry, relieving the burden of neural learning and enhancing efficiency. In this way, the proposed online dense reconstruction framework is both storage-efficient and time-efficient, which is important for application in large-scale scenes with real-time requirements.
More specifically, as shown in Eq. 1, we decouple the complete representation into a coarse representation and a residual one , which are optimized in reconstruction and pose estimation. In general, is expected to be fast and robust, such as TSDF, providing the coarse geometry seamlessly with the input data. is supposed to be more expressive and capture the fine-grained details based on . The coarse estimation is usually constructed based on explicit representations, and the neural networks are adopted for residual refinement . The coarse estimation and residual refinement are performed alternately for reconstruction and pose estimation.
@@
The overview of our method is demonstrated in Figure 2. Given the successive RGB-D inputs, we first leverage the scalable randomized optimization with limited and fixed footprints for the coarse pose estimation (Fig. 2(a)). Keyframes are selected from these input frames. The corresponding initial camera poses are fed into a tiny MLP for the joint residual pose refinement (Fig. 2(b)).
Note that the proposed residual-based BA of the camera poses is optimized by the gradients given by the multi-view constraints based on the reconstructed scene. To improve pose estimation, an explicit-implicit scene representation mixture is introduced simultaneously. The online learning (Fig. 2(c)) of this representation depends on the optimized camera poses. Specifically, we perform ray casting and sample many points along the ray directions accordingly. Embeddings of these points, including the hash and explicit embedding together with positional embedding, are sent to two MLPs for TSDF and RGB residual prediction. The residuals indicating the high-frequency details are added to the explicit coarse base, outputting the final reconstruction. The residual-based BA and residual-based mapping are running alternately with continuously updated global reconstruction and optimized keyframe poses.
@@ 3.2. Residual-based Mapping
In large-scale scenes, various layouts and complex structures make it challenging for the scene representations to remain both expressive and efficient. Shown in Figure 2(c), we employ a mixed representation to address this challenge. This representation consists of an explicit coarse TSDF grid and an implicit neural map , which correspond to and in Eq. 1 respectively. The key insight of the residual-based representations is to obtain the coarse reconstruction in real time and provide the initial surface for the implicit module to efficiently deform and refine. Consequently, given an arbitrary 3D position , the attributes which are the geometry and appearance of can be formed in a residual-based aggregation as:
@@
Given the posed RGB-D frames, the explicit component of the proposed scene representation is constructed via TSDF Fusion (Curless and Levoy, 1996) with a relatively low resolution to store the coarse TSDF and RGB attributes, serving as explicit base of our reconstruction. Similar to some previous implicit-based methods, the implicit component includes two kinds of embedding functions, which are hash embedding (Müller et al., 2022) and positional embedding (Müller et al., 2019), to encode the reconstruction residuals for an arbitrary position upon the . Specifically, adopting this joint embedding in the representation can balance the training efficiency and reconstruction quality (Wang et al., 2023), which is critical for online reconstruction. Furthermore, the residual learning depends on the explicit base, and the final residual-based aggregation to calculate the attributes for a position can be formed based on Eq. 2 as:
@@
Different from the can be reconstructed directly based on the fusion of the posed RGB-D frames, and in our implicit neural map need to be carefully optimized during the reconstruction ( is a constant projection). Therefore, proper supervision for learning and formulating a good loss function are critical here.
Following (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023), we adopt the input RGB-D frames with the estimated poses as the approximation of ground-truth modeling to provide the supervision in a projection-and-measure fashion. Volume rendering is leveraged to output the re-projected RGB-D images based on the weights of points on the casting ray from each image pixel. Similar to (Azinović et al., 2022), the weights are obtained by the Sigmoid functions and as shown in Eq. 5.
Given an image pixel , the TSDF truncation threshold , and the corresponding sampled points on its casting rays with the sampled depth values , the rendered RGB values and depth values can be obtained as Eq. 6 and Eq. 7.
One more thing, due to the resolution of being low, the truncation threshold for it needs to be larger than the one adopted in the residual component . To align these two modules, the explicit base adopted in TSDF component of Eq. 3 needs to be specified as:
The is the clamp function which ensures to be in the range of with threshold . The aligned explicit base is only adopted for the TSDF calculation, while with the direct trilinear interpolation is proper for the RGB prediction.
@@
For the running efficiency, we sample pixels in each iteration for neural learning. More details about sampling strategies can be found in Section 3.4. We use the observed RGB and depth images for supervision. The photometric and geometric losses are as follows:
@@
where is the randomly sampled vertices on the hash grids and means the difference of hash features between the adjacent vertices along these three axes.